### v2

# Unifying and Verifying Mechanistic Interpretations: A Case Study with Group Operations

Abstract
A recent line of work in mechanistic interpretability has focused on reverse-engineering the computation performed by neural networks trained on the binary operation of finite groups. We investigate the internals of one-hidden-layer neural networks trained on this task, revealing previously unidentified structure and producing a more complete description of such models that unifies the explanations of previous works (Chughtai et al., 2023; Stander et al., 2024). Notably, these models approximate equivariance in each input argument. We verify that our explanation applies to a large fraction of networks trained on this task by translating it into a compact proof of model performance, a quantitative evaluation of model understanding. In particular, our explanation yields a guarantee of model accuracy that runs in 30% the time of brute force and gives a $\geq$95accuracy bound for 45% of the models we trained. We were unable to obtain nontrivial non-vacuous accuracy bounds using only explanations from previous works.
1 Introduction
Modern neural network models, despite their widespread deployment and success, remain largely inscrutable in their inner workings, limiting their use in safety-critical settings. The emerging field of mechanistic interpretability seeks to address this issue by reverse engineering the behavior of trained neural networks. One major criticism of this field is the lack of rigorous evaluations of interpretability results; indeed, many works rely on human intuition to determine the quality of an interpretation (Miller, 2019; Casper, 2023; Räuker et al., 2023). This insufficiency of evaluations has proved detrimental to interpretability research: recent work finds many commonly used model interpretations to be imprecise or incomplete (Miller et al., 2024; Friedman et al., 2024).
A simplified research program has focused on toy algorithmic settings, which are made more tractable by the presence of complete mathematical descriptions of the task and dataset (Nanda et al., 2023a; b; Zhong et al., 2024). However, even in these settings, the lack of rigorous evaluations for interpretations is consequential, leading different researchers to come up with divergent explanations for the same empirical phenomena: recently, Chughtai et al. (2023) claimed that models trained on finite groups implement a group composition via representations algorithm, while subsequent work (Stander et al., 2024) studies the same model and task and instead argues that the model implements a coset concentration algorithm.
In this work, we take on the challenge of reconciling their interpretations. We investigate the same setting and find internal model structure that was overlooked by both previous works: the irreducible representations noticed by Chughtai et al. (2023) act by permutation on a discrete set of vectors learned by the model. Based on our observations, we propose a model explanation that unifies those found by both previous works, in particular finding that the model approximates equivariance in each of its input arguments. Following Gross et al. (2024), we then evaluate our interpretation and that of previous work by converting them into compact proofs of lower bounds on model accuracy. Our philosophy is that any rigorous mechanistic knowledge of a model’s inner workings should yield a guarantee on the model’s performance; more complete explanations should yield either tighter performance bounds or a more compact proof, providing a quantitative measure of explanation completeness. We find that our interpretation is indeed an improvement over previous ones because it yields more compact proofs of tigher bounds.
Figure 1: A sketch of the coset and $\rho$-set algorithms in the case of the group $S_{3}$. We visualize elements in $S_{3}$ as points on a hexagon. Top: The coset algorithm introduced in Stander et al. (2024) applies a sieve where different neurons filter out different wrong answers. (Section 4.1). Here, the correct answer is the identity $e$. Bottom: In our proposed algorithm, neurons represent hyperplanes that separate wrong answers from the correct answer. This algorithm is equivariant in each input: in this example, changing the input from $(e,e)$ to $(e,(123))$ transforms the hyperplanes such that the maximal output logit changes from $z=e$ to $z=(123)$ (Section 5).
Our contributions are as follows:
  • We clarify previous mechanistic interpretability results on the group composition task and argue that they provide insufficient evidence to fully explain model behavior (Section 4).
  • We perform our own investigation of the same setting, and provide a more complete explanation (Section 5 and Figure 1). We show that our interpretation unifies the findings of previous works (Section 7).
  • We verify this explanation by translating it into guarantees on model accuracy (Section 6). For a substantial fraction of models, these guarantees are near the true accuracy, providing strong positive evidence for our explanation in these cases (Section 6.2).
2 Related work
Groups, mechanistic interpretability, and grokking
Our work can be seen as a direct follow-up to Chughtai et al. (2023) and Stander et al. (2024), which both perform mechanistic interpretability on one-hidden-layer neural networks trained on the binary operation of finite groups. These papers in turn build on work that studies models trained on modular arithmetic (Nanda et al., 2023a; Zhong et al., 2024), i.e. the binary operation of the cyclic group. Models trained on group composition exhibit the grokking phenomenon (Power et al., 2022), in which a model trained on an algorithmic task generalizes to the test set many epochs after attaining perfect accuracy on the training set. Morwani et al. (2024) study the group composition task from the viewpoint of inductive biases, showing that, for one-hidden-layer models with quadratic activations, the max-margin solution must match the observations of Chughtai et al. (2023).
Evaluation of explanations
Several techniques to evaluate interpretations of models have been suggested, such as causal interventions (Wang et al., 2023) and causal scrubbing (Geiger et al., 2024). We discuss merits and limitations of causal interventions in our setting, which were first explored in Stander et al. (2024). More recently, Gross et al. (2024) use mechanistic interpretability to obtain compact formal proofs of model properties for the max-of-four task. Yip et al. (2024) study the modular arithmetic setting, finding that the ReLU nonlinearities can be thought of as performing numerical integration, and use this insight to compute bounds on model error in linear time.
Equivariance
We find that neural networks trained on group composition learn to be equivariant in both input arguments, i.e. bi-equivariance, despite this condition not being enforced in the architecture. Learned equivariance has been noticed and measured in other settings (Lenc & Vedaldi, 2019; Olah et al., 2020; Gruver et al., 2023). This is distinct from the area of equivariant networks, in which equivariance is enforced by model architecture (Bronstein et al., 2021).
3 Preliminaries
3.1 Mathematical background: groups, actions, representations, $\rho$-sets
This paper uses ideas from finite group theory and representation theory. We provide a rapid and informal introduction to the most important definitions and refer the reader to Section 3 and Appendices D, E, F of Stander et al. (2024) and/or relevant textbooks (Dummit & Foote, 2004; Fulton & Harris, 1991) for more details.
Groups and permutations
A group $G$ is a set with an associative binary operation $\star$ and an identity element $e$ such that every element has an inverse. We write $S_{n}$ for the group of permutations on $n$ elements. Maps [^1] By “maps” we mean group homomorphisms. $G\to S_{n}$ are called permutation representations and are equivalent to actions of $G$ on sets of size $n$. Recall that each permutation $\sigma\in S_{n}$ can be represented as a permutation matrix in ${\mathbb{R}}^{n\times n}$ that applies the permutation $\sigma$ to the basis vectors of ${\mathbb{R}}^{n}$. Thus, any permutation representation of $G$ is a linear representation, i.e. a mapping from $G$ to the group of invertible $n\times n$ matrices $\operatorname{GL}(n,{\mathbb{R}})$. The group operation translates to matrix multiplication.
Irreps
Linear representations that cannot be decomposed into a direct sum of representations of strictly smaller dimension are called irreducible representations or irreps for short. A representation $\rho$ is an irrep if and only if there is no nontrivial subspace that is closed under $\rho(g)$ for all $g\in G$. Every finite group $G$ has a finite set of irreps (up to isomorphism), which we denote by $\operatorname{Irrep}(G)$. [^2] In this paper, we consider irreps over ${\mathbb{R}}$. In particular, all irreps of $S_{n}$ are real. For a discussion of preliminary results for groups with complex irreps, see Appendix I. Any linear representation can be decomposed uniquely into irreps.
$\rho$-sets
By the preceding discussion, given a permutation representation $\tilde{\rho}\colon G\to S_{n}$, we can consider its decomposition into irreps. Let $\rho\in\operatorname{Irrep}(G)$ be one irrep present in this decomposition, acting on some subspace $W\subseteq{\mathbb{R}}^{n}$. Since, by definition, $\tilde{\rho}$ acts on standard basis vectors ${\bm{e}}_{1},\ldots,{\bm{e}}_{n}$ by permutation, the constituent irrep $\rho$ acts on the projection of the basis vectors onto $W$ by the same permutation. We refer to any subset of $W$ that $\rho$ acts on by permutation as a $\rho$-set; in particular, projections of the basis vectors onto $W$ fit this criterion. [^3] The term “$\rho$-set” is our own. All other definitions and notations introduced in this section are standard.
Example: $S_{5}$
Our primary example throughout the main text is the permutation group $S_{5}$. The identity map $S_{5}\to S_{5}$ is a permutation representation. As a linear representation, it is not irreducible, as it fixes the all ones vector $\mathbf{1}\in{\mathbb{R}}^{5}$. Projecting out this vector results in what is called the standard four-dimensional irrep of $S_{5}$; call it $\rho$. This irrep acts by permutation on the projections of the standard basis vectors, so these five vectors in ${\mathbb{R}}^{4}$ form a $\rho$-set. In this example, the $\rho$-set consists of five evenly spaced vectors on the surface of the sphere in $\mathbb{R}^{4}$.
3.2 Task description and model architecture
Our task and architecture are identical to that of previous works (Chughtai et al., 2023; Stander et al., 2024). Fix a finite group $G$. We train a model on the supervised task $\star:G\times G\to G$; i.e., given $x,y\in G$ the task is to predict the product $x\star y\in G$.
We train a one-hidden-layer two-input neural network on this task. The input to the model is two elements $x,y\in G$, embedded as vectors ${\bm{E}}_{l}(x),{\bm{E}}_{r}(y)\in\mathbb{R}^{m}$, which we refer to as the left and right embeddings, respectively. These are multiplied by the left and right linearities ${\bm{W}}_{l},{\bm{W}}_{r}\in{\mathbb{R}}^{m\times m}$, summed, applied with an elementwise ReLU nonlinearity, and finally multiplied by the unembedding matrix ${\bm{U}}\in{\mathbb{R}}^{\lvert G\rvert\times m}$ and summed with the bias ${\bm{w}}_{b}$.
We can simplify the model’s description by noting that the left embedding ${\bm{E}}_{l}$ and left linearity ${\bm{W}}_{l}$ only occur as a product ${\bm{W}}_{l}{\bm{E}}_{l}$, and likewise for the right embedding and linearity. Also, the product between the unembedding ${\bm{U}}$ and the embedding vectors can be decomposed into a sum over the hidden dimensionality $[m]$. Hence, letting ${\bm{w}}_{l}^{i}(x),{\bm{w}}_{r}^{i}(y),{\bm{w}}_{u}^{i}(z)$ denote the $i$th entries of ${\bm{W}}_{l}{\bm{E}}_{l}(x),{\bm{W}}_{r}{\bm{E}}_{r}(y),{\bm{U}}(z)$ respectively, the forward pass is
$f_{\bm{\theta}}(z\mid x,y)={\bm{w}}_{b}(z)+\sum_{i=1}^{m}{\bm{w}}_{u}^{i}(z)\operatorname{ReLU}[{\bm{w}}_{l}^{i}(x)+{\bm{w}}_{r}^{i}(y)],$ (1)
paramaterized by ${\bm{\theta}}=({\bm{w}}_{b},({\bm{w}}_{l}^{i},{\bm{w}}_{r}^{i},{\bm{w}}_{u}^{i})_{i=1}^{m})$. Each vector ${\bm{w}}_{l}^{i},{\bm{w}}_{r}^{i},{\bm{w}}_{u}^{i}$ for $i\in[m]$ can be thought of as a function $G\to{\mathbb{R}}$, and we refer to them as the left, right, and unembedding neurons, respectively. The refer to $i\in[m]$ as the neuron index.
4 Revisiting previous explanations: Cosets and irreps
In this section we recall the coset algorithm described in Stander et al. (2024) and the notion of irrep sparsity in Chughtai et al. (2023). We find that both works correctly identify properties of individual neuron weights, but lack a precise picture of how these neurons combine to compute the group operation.
4.1 Coset concentration
Recall the model architecture Eq 1. For each neuron $i$, the left embeddings are constant on the right cosets of a certain subgroup of $G$, while the right embeddings are constant on the left cosets of a conjugate subgroup. [^4] That is, they are coset concentrated; see Definition 7.2. There is a subset of $G$ common to these two families of left/right cosets, which we call $X_{i}\subseteq G$. On input $x,y\in G$, the left and right embeddings of $i$ sum to zero precisely when $x\star y\in X_{i}$. Meanwhile, the unembedding takes smaller values on elements of $X_{i}$; thus, when $x\star y\not\in X_{i}$, the model’s confidence in $G\setminus X_{i}$ is increased. In the example of $S_{5}$, these $X_{i}$ typically take the form of $X_{ab}:=\{\sigma\in S_{5}\mid\sigma(a)=b\}$. See Stander et al. (2024, Section 5) and Appendix A for more details and Figure 1 (top) for an illustration.
This explanation leaves several things unclear:
  • Even given coset concentration, there are many choices of left/right embeddings that sum to zero whenever $x\star y\in X_{i}$. The choice matters: for example, if there are many input pairs where $x\star y\not\in X_{i}$ but the sum of embeddings is near zero, then the model cannot be expected to perform well. How do we know whether the model has made a good choice?
  • The unembedding is similarly underdetermined: there are many degrees of freedom in choosing weights satisfying the sole constraint of being smaller on $X_{i}$ than on $G\setminus X_{i}$.
  • The bias term is not mentioned, despite being present in the models under study.
In Proposition A.2, we provide a more precise version of this explanation, assuming the model weights satisfy stronger constraints. However, the actual models we investigate do not match these assumptions closely, which is reflected by our failure to convert this explanation into non-vacuous bounds in Appendix C.
4.2 Irrep sparsity
Chughtai et al. (2023) notice that each neuron in the trained model is in the linear span of matrix elements of some irrep $\rho\in\operatorname{Irrep}(G)$; we refer to this condition as irrep sparsity (see Definition 7.1). Based on this observation, they propose that the model:
  1. 1.
    Embeds the input pair $x,y\in G$ as $d\times d$ irrep matrices $\rho(x)$ and $\rho(y)$
  2. 2.
    Uses ReLU nonlinearities to compute the matrix multiplication $\rho(x)\rho(y)=\rho(x\star y)$.
  3. 3.
    Uses ReLU nonlinearities and $\rho(x\star y)$ from the previous step to compute $\operatorname{tr}(\rho(x\star y\star z^{-1}))$, which is maximized at $x\star y\star z^{-1}=e\iff z=x\star y$.
However, because they leave the ReLU computations as a black box, they are unable to fully explain the model’s implemented algorithm. By more carefully investigating which linear combinations of $\rho$ each neuron uses, we deduce a more complete description.
5 Our explanation: Group composition by $\rho$-sets
In this section we explain our findings and suggest a more complete reconstruction of the model’s computation. Crucially, we find that the model approximates a function that is equivariant in each of its arguments; that is, varying the function’s inputs changes the output logits by permuting its values. See Figure 1 (bottom) for a visualization.
5.1 The $\rho$-set circuit
Recall from Eq 1 that $f_{\bm{\theta}}$ can be written as the sum of the contributions of each neuron plus the bias. As seen in Section 4.2, each neuron is in the span of a single irrep $\rho\in\operatorname{Irrep}(G)$. We find that, furthermore, these neurons are actions of $\rho$ on finite $\rho$-sets $\mathcal{B}\subseteq{\mathbb{R}}^{d}$ projected onto one dimension. Moreover, the unembedding weights of each neuron are related to the left and right embedding weights, so that, for some ${\bm{b}},{\bm{b}}^{\prime}$ from a $\rho$-set $\mathcal{B}$ and some projection vector ${\bm{a}}$, up to scaling,
${\bm{w}}^{i}_{u}(z)=-{\bm{b}}^{\top}\rho(z){\bm{b}}^{\prime},\qquad{\bm{w}}^{i}_{l}(x)={\bm{b}}^{\top}\rho(x){\bm{a}},\qquad{\bm{w}}^{i}_{r}(y)=-{\bm{a}}^{\top}\rho(y){\bm{b}}^{\prime}.$ (2)
Additionally, there is one neuron of this form for each of the $\lvert\mathcal{B}\rvert^{2}$ pairs $({\bm{b}},{\bm{b}}^{\prime})$, [^5] To be more precise, we find that there are often multiple neurons per $({\bm{b}},{\bm{b}}^{\prime})$. However, they are scaled such that the sum of neuron contributions corresponding to each pair is uniform. and ${\bm{a}}$ is constant across all such pairs. See Observation B.1 for a full enumeration of our findings.
Based on these observations, we partition neurons into independent $\rho$-set circuits; each circuit is associated with a $\rho\in\operatorname{Irrep}(G)$ of dimension $d$, a finite $\rho$-set $\mathcal{B}\subseteq{\mathbb{R}}^{d}$ that $\rho$ acts on transitively by permutation, and a constant vector ${\bm{a}}\in{\mathbb{R}}^{d}$. Call this circuit $f_{\rho,\mathcal{B}}$; then,
$f_{\rho,\mathcal{B}}(z\mid x,y)=-\sum_{{\bm{b}},{\bm{b}}^{\prime}\in\mathcal{B}}{\bm{b}}^{\top}\rho(z){\bm{b}}^{\prime}\operatorname{ReLU}[{\bm{b}}^{\top}\rho(x){\bm{a}}-{\bm{a}}^{\top}\rho(y){\bm{b}}^{\prime}]$ (3)
Using the $\rho$-set structure of $\mathcal{B}$, we can change variables via $\tilde{\bm{b}}=\rho(x)^{\top}{\bm{b}}=\rho(x^{-1}){\bm{b}}$ and $\tilde{\bm{b}}^{\prime}=\rho(y){\bm{b}}^{\prime}$, and rewrite the sum as
$f_{\rho,\mathcal{B}}(z\mid x,y)=-\sum_{\tilde{\bm{b}},\tilde{\bm{b}}^{\prime}\in\mathcal{B}}\tilde{\bm{b}}^{\top}\rho(x^{-1}\star z\star y^{-1})\tilde{\bm{b}}^{\prime}\operatorname{ReLU}[{\bm{a}}^{\top}(\tilde{\bm{b}}-\tilde{\bm{b}}^{\prime})].$ (4)
Thus $f_{\rho,\mathcal{B}}(z\mid x,y)$ depends only on $x^{-1}\star z\star y^{-1}$; we call such functions bi-equivariant. [^6] To see the equivariance, note that if $f_{\rho,\mathcal{B}}$ is of this form, then, for $g,h\in G$, we have $f_{\rho,\mathcal{B}}(z\mid g\star x,y\star h)=f_{\rho,\mathcal{B}}(g^{-1}\star z\star h^{-1}\mid x,y)$. Since $f_{\bm{\theta}}$ is a sum of such circuits, it is also bi-equivariant. Notice that $z=x\star y$ if and only if $x^{-1}\star z\star y^{-1}=e$, which implies $\rho(x^{-1}\star z\star y^{-1})={\bm{I}}$ [^7] This implication goes both ways if $\rho$ is faithful, which is the case for all irreps of $S_{5}$ with dimension $>1$. Further, terms of the summation in Eq 4 are zero whenever $\tilde{\bm{b}}=\tilde{\bm{b}}^{\prime}$, which is precisely when $\tilde{\bm{b}}^{\top}\rho(e)\tilde{\bm{b}}^{\prime}=\langle\tilde{\bm{b}},\tilde{\bm{b}}^{\prime}\rangle$ is largest. Heuristically, these properties explain why $f_{\rho,\mathcal{B}}(z\mid x,y)$ is maximized when $z=x\star y$. For certain $\rho\in\operatorname{Irrep}(G)$, we can rigorously show that the function is indeed maximized at $z=x\star y$ for any value of ${\bm{a}}$; see Lemma F.3 for details.
Another way to express Eq 4 is as
$f_{\rho,\mathcal{B}}(z\mid x,y)=-\left\langle\rho(x^{-1}\star z\star y^{-1}),\sum_{\tilde{\bm{b}},\tilde{\bm{b}}^{\prime}\in\mathcal{B}}\operatorname{ReLU}[{\bm{a}}^{\top}(\tilde{\bm{b}}-\tilde{\bm{b}}^{\prime})]\tilde{\bm{b}}^{\prime}\tilde{\bm{b}}^{\top}\right\rangle,$ (5)
[…]
5.2 The sign circuit
For irreps of dimension strictly greater than one, we observe that the model learns circuits closely approximating what we describe in Section 5.1. However, for the sign irrep, [^9] The sign irrep maps permutations to ±$\pm 1$ depending on whether they decompose into an odd or even number of transpositions. It is the only nontrivial one-dimensional irrep of $S_{n}$. In general, any real-valued one-dimensional irrep of any group must take values ±$\pm 1$, and the same circuit as described here works. the model is able to use one-dimensionality to avoid the expense of the double summation in Eq 3.
Explicitly, up to scaling, the sign circuit is
$f_{\operatorname{sgn}}(z\mid x,y)=\rho(z)-\rho(z)\operatorname{ReLU}[\rho(x)-\rho(y)]-\rho(z)\operatorname{ReLU}[\rho(y)-\rho(x)]=\rho(x^{-1}\star z\star y^{-1}),$
where $\rho$ is the sign irrep. The circuit comprises two neurons ($-\rho(z)\operatorname{ReLU}[\rho(x)-\rho(y)]$ and $-\rho(z)\operatorname{ReLU}[\rho(y)-\rho(x)]$), and uses the unembedding bias to strip out an extraneous $\rho(z)$ term. See Appendix B.1 for further discussion.
6 Explanations as compact proofs of model performance
Following Gross et al. (2024), we evaluate the completeness of a mechanistic explanation by translating it into a compact proof of model performance. Intuitively, given the model weights and an interpretation of the model, we aim to leverage the interpretation to efficiently compute a guarantee on the model’s global accuracy. More precisely, we construct a verifier program [^10] Formally, a Turing machine. We elide any implementation details related to encoding finite group-theoretic objects as strings, error from finite floating-point precision, etc. $V$ that takes as input the model parameters ${\bm{\theta}}$, the group $G$, and an encoding of the interpretation into a string $\pi$, and returns a real number. We require that $V$ always provides valid lower bounds on accuracy; that is, $V({\bm{\theta}},G,\pi)\leq\operatorname{\alpha_{G}}({\bm{\theta}})$ regardless of the given interpretation $\pi$.  [^11] This soundness requirement prevents $\pi$ from (for example) simply providing the true accuracy $\operatorname{\alpha_{G}}({\bm{\theta}})$ to the verifier. If $V$ takes $\pi$’s veracity for granted, it is no longer sound—$\pi$ could falsely claim the accuracy to be higher than the truth, causing $V({\bm{\theta}},G,\pi)>\operatorname{\alpha_{G}}({\bm{\theta}})$. Our measure of $\pi$’s faithfulness is the tightness of the output guarantee $V({\bm{\theta}},G,\pi)$, i.e. how close it is true to the true accuracy $\operatorname{\alpha_{G}}({\bm{\theta}})$.
Two simple examples of verifiers are:
  1. 1.
    The vacuous verifier $V_{\mathrm{vac}}({\bm{\theta}},G,\emptyset)=0$. This is a valid verifier because 00 is a lower bound on any model’s accuracy.
  2. 2.
    The brute force verifier $V_{\mathrm{brute}}({\bm{\theta}},G,\emptyset)=\operatorname{\alpha_{G}}({\bm{\theta}})$, which takes the model’s weights and runs its forward pass on every input to compute the global accuracy.
While the brute force approach attains the optimal bound by recovering the true accuracy, it is computationally expensive. (Indeed, it is intractable in any real-world setting, where the input space is too large to enumerate.) Notice also that neither example is provided an interpretation $\pi$; without any information about ${\bm{\theta}}$, we cannot expect the verifier to do better than these trivial examples. We aim to construct verifiers that, when $\pi$ is a meaningful interpretation, (1) give non-vacuous guarantees on model accuracy and (2) are compact, i.e. more time-efficient than brute force.
A good understanding of the model’s internals should allow us to compress its description and therefore make reasonable estimates on its error. A more complete explanation should yield a tighter bound, while also reducing the computation cost. Therefore if we think of bounding the accuracy as a trade-off between accuracy and computational cost, a good explanation of the model should push the Pareto frontier outward. See Gross et al. (2024) for a more thorough discussion.
6.1 Constructing compact proofs
The verifier lower-bounds the model’s accuracy by trying to prove for each $x,y\in G$ that the model’s output is maximized at $x\star y$, i.e. that $f_{\bm{\theta}}(x\star y\mid x,y)>\max_{z\neq x\star y}f_{\bm{\theta}}(z\mid x,y)$. Its strategy is:
  1. (1)
    Given model parameters ${\bm{\theta}}$, use the interpretation $\pi$ to construct an idealized model $\tilde{\bm{\theta}}$.
  2. (2)
    For $x,y\in G$, lower bound the margin $f_{\tilde{\bm{\theta}}}(x\star y\mid x,y)-\max_{z\neq x\star y}f_{\tilde{\bm{\theta}}}(z\mid x,y)$.
  3. (3)
    For $x,y\in G$, upper bound the maximum logit distance
    $\max_{z\neq x\star y}\lvert f_{\tilde{\bm{\theta}}}(z\mid x,y)-f_{{\bm{\theta}}}(z\mid x,y)\rvert+f_{\tilde{\bm{\theta}}}(x\star y\mid x,y)-f_{{\bm{\theta}}}(x\star y\mid x,y).$
  4. (4)
    The accuracy lower bound is the proportion of inputs $x,y\in G$ such that the margin lower bound exceeds the distance upper bound. For such input pairs, the margin by which the idealized model’s logit value on the correct answer exceeds the logit value of any incorrect answer is larger than the error between the original and idealized model, so the original model’s logit output must be maximized at the correct answer as well. See Figure 2.
Recall our model architecture Eq 1. The brute-force verifier runs a forward pass with time complexity $O(m\lvert G\rvert)$ over $\lvert G\rvert^{2}$ input pairs, and so takes time $O(m\lvert G\rvert^{3})$ total.  [^12] For simplicity of presentation, we assume all matrix multiplication is performed with the naïve algorithm; that is, the time complexity of multiplying two matrices of size $m\times n$ and $n\times k$ is $O(mnk)$. Our verifiers need to be asymptotically faster than this in order to be performing meaningful compression. Naïvely, though, both steps (2) and (3) take time $O(m\lvert G\rvert^{3})$, no better than brute force. However, we can reduce the time complexity of each to $O(m\lvert G\rvert^{2})$: for (2) by exploiting the internal structure of the idealized model, and for (3) by using Lemma F.1.
Figure 2: Margin lower bound vs. logit distance upper bound over $x,y\in S_{5}$ for $V_{\mathrm{irrep}}$ and $V_{\mathrm{coset}}$ on a single example model. The accuracy lower bound is precisely the number of points for which the margin lower bound is larger than the logit upper bound (shaded region); in this example, the bound from $V_{\mathrm{irrep}}$ is 100% while that from $V_{\mathrm{coset}}$ is 0%. The margin lower bound of $V_{\mathrm{irrep}}$ is constant due to bi-equivariance.
Compact proofs via coset concentration
Intuitively, coset concentration gives a way to perform nontrivial compression—if the interpretation says that a neuron is constant on the cosets of a specific group, then we need only to check one element per coset, instead of a full iteration over $G$. However, the shortcomings listed in Section 4.1 are an obstacle to formalizing this intuition into a compact proof. The verifier $V_{\mathrm{coset}}$ we construct pessimizes over the degrees of freedom not explained by the coset concentration explanation, resulting in accuracy bounds that are vacuous (Section 6.2). See Appendix C for details of $V_{\mathrm{coset}}$’s construction.
Compact proofs via $\rho$-sets
We are able to turn our $\rho$-set circuit interpretation into a proof of model accuracy that gives non-vacuous results on a majority of trained models; see Appendix D for details and Section 6.2 for empirical results.
The interpretation string $\pi$ labels each neuron with its corresponding irrep $\rho$ and its $\rho$-set. The verifier $V_{\mathrm{irrep}}$ is then able to use this interpretation string to construct an idealized version of the input model that implements $\rho$-set circuits (Eq 3) exactly. By bi-equivariance, this idealized model’s accuracy can then be checked with a single forward pass. Finally, $V_{\mathrm{irrep}}$ bounds the distance between the original and the idealized models using Lemma F.1.
6.2 Empirical results for compact proofs
We train 100 one-hidden-layer neural network models from random initialization on the group $S_{5}$. We then compute lower bounds on accuracy obtained by brute force, the cosets explanation, and the $\rho$-sets explation, which we refer to as $V_{\mathrm{brute}},V_{\mathrm{coset}},V_{\mathrm{irrep}}$ respectively. We evaluate these lower bounds on both their runtime [^13] We use the verifier’s time elapsed as a proxy for FLOPs, which is a non-asymptotic measure of runtime. (compactness) and by the tightness of the bound. See Appendix H for full experiment details.
As expected, $V_{\mathrm{brute}}$ obtains the best accuracy bounds (indeed, they are exact), but has a slower runtime than $V_{\mathrm{irrep}}$; see Figure 3. On the other hand, across all experiments, $V_{\mathrm{coset}}$ failed to yield non-vacuous accuracy bounds; while the margin lower bounds of $V_{\mathrm{coset}}$’s idealized models are nonzero, they are swamped by the upper bound on logit distance to the original model; see Figure 2.
Looking again at Figure 3, we see that the accuracy bound due to $\rho$-sets is bimodal: the verifier $V_{\mathrm{irrep}}$ obtains a bound of near 100% for roughly half the models, and a vacuous bound of 0% for another half. Investigating the models for which $V_{\mathrm{irrep}}$ does not obtain a good bound, we are able to discover for many of them aspects in which they deviate from our $\rho$-sets explanation:
  • The ${\bm{a}}$ projection vector (Eq 3) fails to be constant across terms of the double sum over ${\bm{b}},{\bm{b}}^{\prime}\in\mathcal{B}$, so the change of variables showing bi-equivariance (Eq 4) is invalid. [^14] If ${\bm{a}}$ depends on $({\bm{b}},{\bm{b}}^{\prime})$, then there is a remaining dependence on $(\rho(x)\tilde{\bm{b}},\rho(y^{-1})\tilde{\bm{b}}^{\prime})$ inside the ReLU after changing variables. We find that such models have poorer cross-entropy loss and larger weight norm than those with constant ${\bm{a}}$, suggesting they have converged to a suboptimal local minimum (Figure 4 in Appendix 4).
  • The double sum over $\mathcal{B}$ (Eq 3) misses some ${\bm{b}},{\bm{b}}^{\prime}$ pairs. Again, we are unable to prove bi-equivariance when this happens. We speculate that in this case the model is approximating the discrete summation by a numerical integral à la Yip et al. (2024).
Although these cases are failures of our $\rho$-sets interpretation to explain the model, their presence can be seen as a success for compact proofs as a measure of interpretation completeness. For models we do not genuinely understand, we are unable to achieve non-vacuous guarantees on accuracy.
Figure 3: Accuracy bound vs. computation time for $V_{\mathrm{irrep}}$ and $V_{\mathrm{brute}}$ on 100 models trained on $S_{5}$. Points in green ($V_{\mathrm{irrep}}$ unexpl) are models for which we find by inspection that our $\rho$-sets explanation does not hold (See Section 6.2). Mean accuracy bound is 100% for $V_{\mathrm{brute}}$, 0% for $V_{\mathrm{coset}}$ (not shown), 51.0% for $V_{\mathrm{irrep}}$, and 82.3% for $V_{\mathrm{irrep}}$ when excluding models known not to fit our $\rho$-sets explanation (38% of total). Proportion of models with $V_{\mathrm{irrep}}$ accuracy bound $\geq 95\%$ is 45.0% for all models and 72.6% excluding unexplained. Mean time elapsed is 2.01s for $V_{\mathrm{brute}}$ and 0.61s for $V_{\mathrm{irrep}}$. The asymptotic time complexity of $V_{\mathrm{brute}}$ is $O(m\lvert G\rvert^{3})$ while that of $V_{\mathrm{irrep}}$ is $O(m\lvert G\rvert^{2})$.
7 Discussion
7.1 Relating irrep sparsity, coset concentration, and $\rho$-sets
This paper and prior works observe multiple properties of neurons viewed as functions $G\to{\mathbb{R}}$:
Definition 7.1 (Chughtai et al. 2023).
A function $f\colon G\to{\mathbb{R}}$ is irrep sparse if it is a linear combination of the matrix entries of an irrep of $G$. That is, there exists a $\rho\in\operatorname{Irrep}(G)$ of dimension $d$ and a matrix ${\bm{A}}\in{\mathbb{R}}^{d\times d}$ such that $f(g)=\operatorname{tr}(\rho(g){\bm{A}})$.
Definition 7.2 (Stander et al. 2024).
A function $f\colon G\to{\mathbb{R}}$ is coset concentrated if there exists a nontrivial subgroup $H\leq G$ such that $f$ is constant on the cosets (either left or right) of $G$.
Definition 7.3.
A function $f\colon G\to{\mathbb{R}}$ is a projected $\rho$-set for $\rho\in\operatorname{Irrep}(G)$ of dimension $d$ if there exist ${\bm{a}},{\bm{b}}\in{\mathbb{R}}^{d}$ such that $f(g)={\bm{a}}^{\top}\rho(g){\bm{b}}$ and ${\bm{b}}$ is in a $\rho$-set with nontrivial stabilizer.
In this section, we clarify the logical relationships between these three properties.
Projected $\rho$-sets are irrep-sparse
This fact, while immediate from definitions, resolves a mystery from Chughtai et al. (2023, Figure 7): why is the standard irrep $\rho_{\mathrm{std}}$ of $S_{5}$ learned significantly more frequently than the sign-standard irrep $\rho_{\mathrm{sgnstd}}$, when both have the same dimensionality $\dim\rho_{\mathrm{std}}=\dim\rho_{\mathrm{sgnstd}}=4$? The answer is that the smallest $\rho_{\mathrm{std}}$-set has size 5, while the smallest $\rho_{\mathrm{sgnstd}}$-set has size 10 (Appendix G). Thus a minimal complete $\rho_{\mathrm{std}}$-set circuit needs $255^{2}=$ 25 neurons, while a minimal complete $\rho_{\mathrm{sgnstd}}$-set circuit needs $10010^{2}=$ 100. The frequency with which $\rho\in\operatorname{Irrep}(G)$ is learned correlates not with its dimensionality, but with the size of the minimal $\rho$-set.
Projected $\rho$-sets are coset concentrated
Our $\rho$-set interpretation immediately explains coset concentration: the map $g\mapsto\rho(g){\bm{b}}$ is constant on precisely the cosets of the stabilizer $\operatorname{Stab}_{G}({\bm{b}}):=\{g\in G\mid\rho(g){\bm{b}}={\bm{b}}\}$. Further, since the action of $G$ is transitive, $\operatorname{Stab}_{G}({\bm{b}}^{\prime})$ for any other element ${\bm{b}}^{\prime}$ is a conjugate of $\operatorname{Stab}_{G}({\bm{b}})$; as a consequence left and right preactivations of each neuron concentrate on cosets of conjugate subgroups.
Coset concentration fails to explain irrep sparsity
Stander et al. (2024) partially explain irrep sparsity via coset concentration. We paraphrase their key lemma here:
Lemma 7.4 (Stander et al., 2024).
Let $H$ be a subgroup of $G$. The Fourier transform of a function constant on the cosets of $H$ is nonzero only at the irreducible components of the permutation representation corresponding to the action of $G$ on $G/H$.
This lemma fails to fully explain irrep sparsity, since the permutation representation of $G$ on $G/H$ is never itself an irrep; it always contains the trivial irrep, possibly among others. Thus, it does not explain why the models’ neurons are supported purely on single irreps and not, say, on linear combinations of each irrep with the trivial irrep.
Notice also that, since there are many more subgroups than irreps (Stander et al., 2024, Appendix G.3), most subgroups $H\leq G$ have corresponding actions of $G$ on $G/H$ that decompose into more than two irreps; otherwise, since the decomposition is unique, there would be at most $\lvert\operatorname{Irrep}(G)\rvert^{2}$ total possibilities. Explicit examples of subgroups whose indicators are supported on more than two irreps are given in Appendix G.
Projected $\rho$-sets are equivalent to the conjunction of irrep sparsity and coset concentration
Note that neither irrep sparsity nor coset concentration alone is equivalent to the condition of being a projected $\rho$-set. For an example of an irrep-sparse function that is not a $\rho$-set, consider the function $\chi(g)=\operatorname{tr}(\rho(g))$. If $\rho$ is faithful and has nontrivial stabilizer (for example, the standard 4-dimensional irrep of $S_{5}$), then $\chi$ cannot be a projected $\rho$-set, because it is maximized uniquely at the identity (Chughtai et al., 2023, Theorem D.7) and thus is not coset-concentrated. To see that coset concentration does not imply being a projected $\rho$-set, recall from above that there are coset-concentrated functions that are not irrep-sparse, but all projected $\rho$-sets are irrep sparse.
On the other hand, if $f\colon G\to{\mathbb{R}}$ is both irrep-sparse and coset-concentrated, then it must be a projected $\rho$-set; for a proof see Lemma F.4. Hence, if we consider by itself a single embedding of a single neuron on $f_{\bm{\theta}}$, then our Observation B.2(2) is logically equivalent to the combination of observations from previous works. However, our perspective gives us more insight into the relationship between left/right embedding and unembedding neurons (Observation B.2(1)) as well as the relationship between different neurons (Observation B.2(6,7)).
7.2 Evaluation of explanations
Limitations of the $\rho$-sets interpretation
The $\rho$-set explanation we provide has a rather limited scope: we only claim to understand roughly half of the models we examine, all of which are trained on $S_{5}$. On the other hand, all of the models we examine satisfy both irrep sparsity and coset concentration. It was by trying and failing to obtain non-vacuous compact guarantees on model accuracy that we discovered that our understanding of some models is still incomplete. Hence, compact proofs are a quantitative means of detecting gaps in proposed model explanations.
Causal interventions
In an attempt to verify the validity of the coset concentration interpretation, Stander et al. (2024) perform a series of causal interventions. While the results do not yield evidence that their interpretation is incorrect, they also do not provide strong evidence that it is correct. Indeed, we perform the same interventions on a model for which we know the cosets explanation cannot hold, and find results in the same direction; see Appendix E. Therefore, causal interventions can yield strong negative evidence against an explanation, but provide weaker positive evidence.
Compact proofs
For the models we do understand, we obtain tight accuracy bounds in significantly less time than brute force, providing strong positive evidence for our understanding. Hence, we view the compact proof approach as complementary to the causal intervention one: A tight bound is strong positive evidence for an explanation, whereas a poor or vacuous bound does not give us strong negative evidence. Indeed, the coset interpretation does make nontrivial observations and yield partial explanations of model performance, but this is not reflected in the vacuous bounds we obtain. It may be the case that some of the pessimizations we use in our construction of $V_{\mathrm{coset}}$ were unnecessarily strong; the translation from an informal explanation into a rigorous bound is itself informal, by necessity.
8 Conclusion
Multiple previous works (Chughtai et al., 2023; Stander et al., 2024) have examined the group composition setting and claimed a mechanistic understanding of model internals. However, as we have demonstrated here, these works left much internal structure unrevealed. Our own interpretation incorporates this structure, resulting in a more complete explanation that unifies previous attempts.
We verify our explanation with compact proofs of model performance, and obtain strong positive evidence that it holds for a large fraction of the models we investigate. For models where we fail to obtain bounds, we find that many indeed do not fit our proposed interpretation. Compact proofs thus provide rigorous and quantitative positive evidence for an interpretation’s completeness. We see this work as a preliminary step towards a more rigorous science of interpretability.
Reproducibility Statement
See Appendix H for experiment details. Code for reproducing our experiments can be found at https://anonymous.4open.science/r/groups-E024.
Acknowledgements
We thank Jesse Hoogland, Daniel Filan, Ronak Mehta, Mathijs Henquet, and Alice Rigg for helpful discussions and feedback on drafts of this paper. WW was partially supported by a grant from the Long-Term Future Fund. This research was partly conducted during the ML Alignment & Theory Scholars Program.
Author Contributions
Wilson Wu
Led the engineering and ran the bulk of all experiments. Designed, implemented, and ran compact proofs. Proved majority of the theoretical results. Contributed to writing of paper.
Louis Jaburi
Initiated the project and suggested experiments to run. Contributed to writing of paper.
Jacob Drori
Made contributions leading to discovery of the $\rho$-set circuit: pushed to study the structure of ${\bm{A}}_{i},{\bm{B}}_{i},{\bm{C}}_{i}$ (defined in Observation B.1), predicted Observation 1; predicted $\{{\bm{b}}_{i}\}$ is a $\rho$-set; noticed ${\bm{b}}_{i}$ and ${\bm{c}}_{i}$ vary independently, leading to the double-sum in Equation 3.
Jason Gross
Advised the project, including developing proof strategies and providing feedback on experimental results and presentation.
Appendix A Coset concentration in detail
The main observation of Stander et al. (2024) is that neurons are concentrated on the cosets of subgroups. More precisely,
Observation A.1 (Stander et al., 2024).
For each $i\in[m]$, there exists a subgroup $H_{i}\leq G$ and element $g_{i}\in G$ such that the left and right neurons are approximately constant on the right cosets of $H_{i}$ and the left cosets of the conjugate subgroup $K_{i}=g_{i}H_{i}g_{i}^{-1}$, [^15] Throughout the appendix, we denote the group multiplication of $x,y\in G$ by $xy$ instead of by $x\star y$. respectively:
$xy^{-1}\in H_{i}\implies{\bm{w}}_{l}^{i}(x)\approx{\bm{w}}_{l}^{i}(y),\quad\quad x^{-1}y\in K_{i}^{-1}\implies{\bm{w}}_{r}^{i}(x)\approx{\bm{w}}_{r}^{i}(y).$ (6)
This forces the summed embeddings ${\bm{w}}_{l}^{i}(x)+{\bm{w}}_{r}^{i}(y)$ to be approximately constant on the double cosets $H_{i}\backslash G/K_{i}$. Further, it is observed that
${\bm{w}}_{l}^{i}(x)+{\bm{w}}_{r}^{i}(y)\approx 0\iff xy\in H_{i}xyg_{i}K_{i}=H_{i}g_{i}^{-1}.$ (7)
If we include several additional assumptions, this observation is sufficient for $f_{\bm{\theta}}$ to attain perfect accuracy. Note it was not explicitly investigated in Stander et al. (2024) to what extent these additional conditions are met.
Proposition A.2.
Suppose the model parameters ${\bm{\theta}}$ are such that Observation A.1 holds, the unembeddings satisfy
$\max_{z\not\in H_{i}g_{i}^{-1}}{\bm{w}}_{u}^{i}(z)=\min_{z\not\in H_{i}g_{i}^{-1}}{\bm{w}}_{u}^{i}(z)>\max_{z\in H_{i}g_{i}^{-1}}{\bm{w}}_{u}^{i}(z)$ (8)
and the bias ${\bm{w}}_{b}$ is zero. Further, defining
$s{(H,g)}=\min_{\begin{subarray}{c}x,y\in G\\ xy\not\in H_{i}g_{i}^{-1}\end{subarray}}\sum_{\begin{subarray}{c}i\in[m]\\ (H_{i},g_{i})=(H,g)\end{subarray}}\operatorname{ReLU}[{\bm{w}}_{l}^{i}(x)+{\bm{w}}_{r}^{i}(y)],$ (9)
suppose every singleton $\{z\}$ for $z\in G$ can be written as an intersection of sets from the family
$\{G-H_{i}g_{i}^{-1}\mid i\in[m],s{(H_{i},g_{i})}>0\}$
Then, $\operatorname{\alpha_{G}}(f)=1$.
Proof.
Let $x,y,z\in G$ with $z\neq xy$. Then,
$f(xy\mid x,y)-f(z\mid x,y)$ $=\sum_{i=1}^{m}({\bm{w}}^{i}_{u}(xy)-{\bm{w}}^{i}_{u}(z))\operatorname{ReLU}[{\bm{w}}_{l}^{i}(x)+{\bm{w}}_{r}^{i}(y)]$
$\geq\sum_{(H,g)}s{(H,g)}\mathbf{1}\{xy\not\in H_{i}g_{i}^{-1}\}\sum_{\begin{subarray}{c}i\in[m]\\ (H_{i},g_{i})=(H,g)\end{subarray}}({\bm{w}}^{i}_{u}(xy)-{\bm{w}}^{i}_{u}(z))$
$\geq 0,$
To see that the inequality is strict, choose $i$ such that $s{(H_{i},g_{i})}>0$ with $z\in H_{i}g_{i}^{-1}$ and $xy\not\in H_{i}g_{i}^{-1}$. ∎
Appendix B $\rho$-set circuits in detail
Chughtai et al. (2023) observe that neurons are irrep-sparse. That is,
Observation B.1 (Chughtai et al., 2023).
For each $i\in[m]$, there exists a real-valued irrep ${\rho_{i}\colon G\to\operatorname{GL}(d_{i},{\mathbb{R}})}$ of degree $d_{i}$ as well as ${\bm{A}}_{i},{\bm{B}}_{i},{\bm{C}}_{i}\in{\mathbb{R}}^{d_{i}\times d_{i}}$ such that
${\bm{w}}_{l}^{i}(x)\approx\operatorname{tr}(\rho(x){\bm{A}}_{i}),\quad{\bm{w}}_{r}^{i}(x)\approx\operatorname{tr}(\rho(x){\bm{B}}_{i}),\quad{\bm{w}}_{u}^{i}(x)\approx\operatorname{tr}(\rho(x){\bm{C}}_{i}).$
However, this is not sufficient to fully describe how the model computes the group operation; in particular the ReLU nonlinearities are left as black boxes.
We observe further structure:
Observation B.2 (Ours).
Let $(\rho_{i},{\bm{A}}_{i},{\bm{B}}_{i},{\bm{C}}_{i})_{i=1}^{m}$ be as in Observation B.1. Suppose that all irreps of $G$ over ${\mathbb{C}}$ are real-valued; in particular, this is the case for $S_{n}$ [^16] Equivalently, the Frobenius-Schur indicator of each $\rho\in\operatorname{Irrep}(G)$ is positive. We briefly consider complex and quaternionic irreps in Appendix I.
  1. 1.
    ${\bm{C}}_{i}\approx r_{i}{\bm{B}}_{i}{\bm{A}}_{i}$ for some $r_{i}>0$.
  2. 2.
    Each of ${\bm{A}}_{i},{\bm{B}}_{i},{\bm{C}}_{i}$ are rank one. Hence, we may write
    ${\bm{A}}_{i}\approx s_{i}{\bm{a}}_{i}{\bm{b}}_{i}^{\top},\quad{\bm{B}}_{i}\approx t_{i}{\bm{c}}_{i}{\bm{d}}_{i}^{\top},\quad{\bm{C}}_{i}\approx r_{i}s_{i}t_{i}\langle{\bm{d}}_{i},{\bm{a}}_{i}\rangle{\bm{c}}_{i}{\bm{b}}_{i}^{\top},$ (10)
    where $s_{i},t_{i}\in{\mathbb{R}}$ and ${\bm{a}}_{i},{\bm{b}}_{i},{\bm{c}}_{i},{\bm{d}}_{i}\in{\mathbb{R}}^{d_{i}}$ are unit vectors.
  3. 3.
    ${\bm{a}}_{i}={\bm{d}}_{i}$.
Further, fix an irrep $\rho$ of $G$ and consider $I_{\rho}=\{i\in[m]\mid\rho_{i}=\rho\}$, the subset of neurons corresponding supported on $\rho$. Then,
  1. 4.
    ${\bm{a}}_{i}$ is approximately constant on $I_{\rho}$. That is, there exists a unit vector ${\bm{a}}_{\rho}\in{\mathbb{R}}^{d_{i}}$ such that
    $\forall i\in I_{\rho}:{\bm{a}}_{i}\approx{\bm{d}}_{i}\approx{\bm{a}}_{\rho}.$
  2. 5.
    $\{{\bm{b}}_{i}\}_{i\in I_{\rho}}\approx\{-{\bm{c}}_{i}\}_{i\in I_{\rho}}$.
  3. 6.
    Each $\{{\bm{b}}_{i}\}_{i\in I_{\rho}}$ can be partitioned into $\{\mathcal{B}_{\rho,q}\}_{q}$ such that, for each neuron $i$, the corresponding vectors ${\bm{b}}_{i}$ and $-{\bm{c}}_{i}$ must belong to the same partition. Further, $\rho$ acts on each partition by permutations; that is, $\rho$ induces a left $G$-set structure on every $\mathcal{B}_{\rho,q}$. We say that such $\mathcal{B}_{\rho,q}$ is a $\rho$-set.
  4. 7.
    If $\dim\rho>1$, then, for each partition $\mathcal{B}_{\rho,q}$, there exists $c_{\rho,q}\in\mathbb{R}$ such that, for every pair ${\bm{b}},{\bm{b}}^{\prime}\in\mathcal{B}_{\rho,q}$,
    $\sum_{i\in I_{\rho}}s_{i}^{2}t_{i}^{2}r_{i}\mathbf{1}\{{\bm{b}}_{i}={\bm{b}},{\bm{c}}_{i}=-{\bm{b}}^{\prime}\}\approx c_{\rho,q}.$
  5. 8.
    If $\dim\rho=1$, then, since we assume $\rho$ is real-valued, it must be either trivial or the sign irrep. We observe that the former never occurs. In the latter case, $\mathcal{B}_{\rho}=\{\pm 1\}$, and there exist $c_{+},c_{-}\in\mathbb{R}$ such that
    $c_{+}$ $\approx\sum_{i\in I_{\rho}}s_{i}^{2}t_{i}^{2}r_{i}\mathbf{1}\{{\bm{b}}_{i}=1,{\bm{b}}_{j}=1\}\approx\sum_{i\in I_{\rho}}s_{i}^{2}t_{i}^{2}r_{i}\mathbf{1}\{{\bm{b}}_{i}=-1,{\bm{b}}_{j}=-1\},$
    $c_{-}$ $\approx\sum_{i\in I_{\rho}}s_{i}^{2}t_{i}^{2}r_{i}\mathbf{1}\{{\bm{b}}_{i}=1,{\bm{b}}_{j}=-1\}\approx\sum_{i\in I_{\rho}}s_{i}^{2}t_{i}^{2}r_{i}\mathbf{1}\{{\bm{b}}_{i}=-1,{\bm{b}}_{j}=1\}.$
  6. 9.
    The bias ${\bm{w}}_{b}$ satisfies
    ${\bm{w}}_{b}(z)\approx(c_{-}-c_{+})\operatorname{sgn}(z).$
Proposition B.3.
If $f_{\bm{\theta}}$ satisfies Observations B.1 and B.2 exactly, then for all $x,y,z,g_{1},g_{2}\in G$,
$f_{\bm{\theta}}(z\mid g_{1}x,yg_{2})=f_{\bm{\theta}}(g_{1}^{-1}zg_{2}^{-1}\mid x,y).$
We say that $f_{\bm{\theta}}$ is bi-equivariant. In particular,
$f_{\bm{\theta}}(z\mid x,y)=f_{\bm{\theta}}(x^{-1}zy^{-1}\mid e,e),$
so such $f_{\bm{\theta}}$ depends only on $x^{-1}zy^{-1}$. Observe that $x^{-1}zy^{-1}=e$ iff $z=xy$.
Proof.
In the notation of Observation B.2, for each $\rho\in\operatorname{Irrep}(G)$ and partition $\mathcal{B}_{\rho,q}$, let $f^{\rho,q}_{\bm{\theta}}=\sum_{i\in I_{\rho}}\mathbf{1}\{{\bm{b}}_{i}\in\mathcal{B}_{\rho,q}\}f^{i}_{\bm{\theta}}$. Then, if $f_{\bm{\theta}}$ satisifes Observation B.2 exactly,
  • If $\dim\rho>1$, for every $\mathcal{B}_{\rho,q}$, there exists ${\bm{a}}\in\mathbb{R}^{\dim\rho},c\in\mathbb{R}$ (all depending on $\rho$) such that, by re-indexing neurons,
    $f_{\bm{\theta}}^{\rho,q}(z\mid x,y)$ $=-c\sum_{i,j=1}^{k}{\bm{b}}_{i}^{\top}\rho(z){\bm{b}}_{j}\operatorname{ReLU}[{\bm{b}}_{i}^{\top}\rho(x){\bm{a}}-{\bm{a}}\rho(y){\bm{b}}_{j}]$
    $=\sum_{i,j=1}^{k}{\bm{b}}_{i}^{\top}\rho(x^{-1}zy^{-1}){\bm{b}}_{j}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]$
    $=\left\langle\rho(x^{-1}zy^{-1}),-c\sum_{i,j}^{k}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{b}}_{j}{\bm{b}}_{i}^{\top}\right\rangle.$
  • If $\dim\rho=1$, then $\rho$ must be the sign irrep. In this case,
    $f_{\bm{\theta}}^{\rho}$ $(z\mid x,y)+{\bm{w}}_{b}(z)$
    $=c_{+}\rho(z)\operatorname{ReLU}[\rho(x)+\rho(y)]+c_{+}\rho(z)\operatorname{ReLU}[-\rho(y)-\rho(x)]-c_{+}\rho(z)$
    $\qquad-c_{-}\rho(z)\operatorname{ReLU}[\rho(x)-\rho(y)]-c_{-}\rho(z)\operatorname{ReLU}[\rho(y)-\rho(x)]+c_{-}\rho(z)$
    $=c_{+}(\rho(z)\lvert\rho(x)+\rho(y)\rvert-\rho(z))-c_{-}(\rho(z)\lvert\rho(x)-\rho(y)\rvert-\rho(z))$
    $=c_{+}(\rho(z)(1+\rho(xy))-\rho(z))-c_{-}(\rho(z)(1-\rho(xy))-\rho(z))$
    $=(c_{+}+c_{-})\rho(zxy)$
    $=(c_{+}+c_{-})\rho(x^{-1}zy^{-1}).$
B.1 Sign irrep and modular arithmetic
Let us briefly extend our attention to groups with complex-valued irreps and consider cyclic groups ${\mathbb{Z}}/p{\mathbb{Z}}$, i.e. arithmetic modulo $p$. All irreps of cyclic groups are one-dimensional, looking like $\rho(x)=e^{2\pi ikx/p}=\cos(2\pi kx/p)+i\sin(2\pi kx/p)$. The modular arithmetic setting is studied by Yip et al. (2024), where it is found that models use an approximation trick involving the sum-of-cosines formula and integration over a single variable:
[…]
$\qquad=\left\lvert\cos\left(\dfrac{x-y}{2}\right)\right\rvert\dfrac{2\cos(x+y-z)}{3}.$
Here, analogously to the sign circuit of our setting, the sum of the group elements’ embeddings is expressed as the embedding of their sum in order to compute the desired inequality with only a single sum/integral, instead of a double sum. (In this case, the extraneous $\lvert\cos((x-y)/2)\rvert$ is not removed; indeed, this additional term is the “Achilles’ heel” of this strategy, and necessitates that the model use multiple irreps, even though each one is faithful (Zhong et al., 2024).)
B.2 Constant projection vectors
Figure 4: Plots of normalized variance $\mathbb{E}_{i}[\lVert{\bm{a}}_{i}-\mathbb{E}_{i}{\bm{a}}_{i}\rVert_{2}^{2}]/\mathbb{E}_{i}[\lVert{\bm{a}}_{i}\rVert_{2}^{2}]$ vs. model loss and weight norm, where ${\bm{a}}_{i}$ is the projection vector for neuron $i$, and expectation is taken across neurons within the 4d standard irrep of $S_{5}$. Each point is one model out of 100 trained on $S_{5}$. Notice that constant ${\bm{a}}_{i}$ across neurons is correlated with better model performance and lower weight norm.
Why does the network learn to set the ${\bm{a}}$ vector constant across neurons? A heuristic explanation is that such a constant vector ${\bm{a}}$ is one way to enforce bi-equivariance, which then leads the margin attained to be uniform across inputs $x,y\in G$. Morwani et al. (2024) show that this uniform margin must necessarily be the case at a maximum margin solution; further, models trained on cross-entropy loss in the zero weight decay limit indeed attain the maximum margin (Wei et al., 2019).
However, we do see that some fraction of trained models do not have constant ${\bm{a}}$; we do not have a full understanding of these models, and thus are unable to non-vacuously bound accuracy. We notice that these models tend to have inferior performance and higher weight norm, suggesting that they have converged to a poor local minimum by chance; see Figure 4. Further, even in these cases, we observe that the ${\bm{a}}_{i}$ all lie in a two-dimensional subspace.
Appendix C Accuracy bounds via coset concentration
We construct a verifier $V_{\mathrm{coset}}$ that takes as input ${\bm{\theta}}$ and an interpretation string
$\pi=((H_{i},g_{i}))_{i=1}^{m},$
and returns a lower bound on $\operatorname{\alpha_{G}}({\bm{\theta}})$ in time $O(m\lvert G\rvert^{2})$.
For each $x,y\in G$, the verifier $V_{\mathrm{coset}}$ computes a lower bound on the margin
$M(z)\leq\min_{\begin{subarray}{c}x,y,z^{\prime}\in G\\ xy=z\neq z^{\prime}\end{subarray}}f_{\bm{\theta}}(xy\mid x,y)-f_{\bm{\theta}}(z^{\prime}\mid x,y)$
by doing the following:
  1. 1.
    Check that the $(H_{i})_{i=1}^{m}$ are each subgroups of $G$ and construct the conjugate subgroups $K_{i}=g_{i}H_{i}g_{i}^{-1}$ in time $O(\sum_{i=1}^{m}\lvert H_{i}\rvert^{2})$.
  2. 2.
    Construct idealized parameters $\tilde{\bm{\theta}}$ by averaging over the cosets given in $\pi$. Explicitly,
    $\tilde{\bm{w}}_{l}^{i}(x)$ $=\lvert H_{i}\rvert^{-1}\sum_{x^{\prime}\in H_{i}x}{\bm{w}}_{l}^{i}(x^{\prime})$
    $\tilde{\bm{w}}_{r}^{i}(y)$ $=\lvert K_{i}\rvert^{-1}\sum_{y^{\prime}\in yK_{i}}{\bm{w}}_{r}^{i}(y^{\prime})$
    $\tilde{\bm{w}}_{u}^{i}(z)$ =$=\begin{cases}\lvert G-H_{i}g_{i}^{-1}\rvert^{-1}\sum_{z^{\prime}\not\in H_{i}g_{i}^{-1}}{\bm{w}}_{u}^{i}(z^{\prime})&z\not\in H_{i}g_{i}\\ \min\{{\bm{w}}_{u}^{i}(z),\min_{z^{\prime}\not\in H_{i}g_{i}}\tilde{\bm{w}}_{u}^{i}(z^{\prime})\}&z\in H_{i}g_{i}\end{cases}$
    $\tilde{\bm{w}}_{b}$ $={\bm{0}}.$
    This takes time $O(\lvert G\rvert m)$.
  3. 3.
    Compute $s{(H_{i},g_{i})}$ for each $i$ according to Eq 9. This takes time $O(\sum_{i=1}^{m}\lvert H_{i}\backslash G\rvert^{2})$.
  4. 4.
    For each $z\in G$, compute
    $M(z)=\min_{\begin{subarray}{c}z^{\prime}\in G\\ z^{\prime}\neq z\end{subarray}}\sum_{(H,g)}s{(H,g)}\mathbf{1}\{z\not\in Hg^{-1}\}\sum_{\begin{subarray}{c}i\in[m]\\ (H_{i},g_{i})=(H,g)\end{subarray}}({\bm{w}}^{i}_{u}(z)-{\bm{w}}^{i}_{u}(z^{\prime}))\$
    This takes time $O(m\lvert G\rvert^{2})$.
  5. 5.
    Lemma F.2 gives an upper bound on the $\ell_{\infty}$ norm between the output logits of original and idealized models for each $x,y\in G$ in time $O(m\lvert G\rvert^{2})$. Sum the difference in logit values of the original and idealized models on the correct answer, which again can be computed in time $O(m\lvert G\rvert^{2})$. This gives
    $L(x,y)\geq\lVert f_{\bm{\theta}}(\cdot\mid x,y)-f_{\tilde{\bm{\theta}}}(\cdot\mid x,y)\rVert_{\infty}+f_{\tilde{\bm{\theta}}}(xy\mid x,y)-f_{{\bm{\theta}}}(xy\mid x,y).$ (11)
    The final accuracy bound is
    $\Pr_{z\sim\operatorname{Unif}(G)}[M(z)>\max_{\begin{subarray}{c}x,y\in G\\ xy=z\end{subarray}}L(x,y)].$
The total time complexity is dominated by the last two steps. Soundness, i.e. $\forall\pi:V_{\mathrm{coset}}({\bm{\theta}},G,\pi)\leq\operatorname{\alpha_{G}}({\bm{\theta}})$, is obvious from construction and Proposition A.2.
The bias term
The description of the coset circuit in Stander et al. (2024) makes no mention of the unembedding bias term, although it is present in the models they train for their experiments. In practice, we find that the bias term is large. The maximum minus minimum value of the bias would then be added directly to $L(x,y)$ for every $x,y\in G$ if we were to run $V_{\mathrm{coset}}$ as written. Thus, in addition, we train models without an explicit bias term for our $V_{\mathrm{coset}}$ experiments. We find that such models are qualitatively similar to models with an explicit bias; the missing bias term is simply added uniformly to each unembedding weight ${\bm{w}}^{i}_{u}$. Perhaps because of this, we are still unable to obtain nonvacuous bounds from $V_{\mathrm{coset}}$ even for these bias-less models.
Appendix D Accuracy bounds via $\rho$-sets
We construct a verifier $V_{\mathrm{irrep}}$ that takes as input ${\bm{\theta}}$ and an interpretation string $\pi$ and returns a lower bound on $\operatorname{\alpha_{G}}({\bm{\theta}})$ in time $O(m\lvert G\rvert^{2})$. The interpretation $\pi$ comprises $((\rho_{i},q_{i},{\bm{a}}_{i},{\bm{b}}_{i},c_{i})_{i=1}^{m})$, where $q_{i}$ is the index of the corresponding $\rho$-set $\mathcal{B}_{\rho,q}=\{{\bm{b}}_{i}\mid\rho_{i}=\rho,q_{i}=q\}$. The verifier then does
  1. 1.
    Check that each $\mathcal{B}_{\rho,q}$ is indeed a $\rho$-set. [^17] Note that this is sufficient for bi-equivariance to go through; we do not need to explicitly check that $\rho\in\operatorname{Irrep}(G)$. This takes time $O(\sum_{\rho,q}\lvert G\rvert\lvert\mathcal{B}_{\rho,q}\rvert^{2}\dim(\rho)^{2})$. Since $m\geq\sum_{\rho,q}\lvert\mathcal{B}_{\rho,q}\rvert^{2}$ and $\sum_{\rho\in\operatorname{Irrep}(G)}\dim(\rho)^{2}=\lvert G\rvert$, this is upper-bounded by $O(\lvert G\rvert^{2}m)$.
  2. 2.
    Within neurons corresponding to each $(\rho,q)$, check that ${\bm{a}}_{i}$ is constant. This again takes time no more than $O(\lvert G\rvert m)$.
  3. 3.
    For each $(\rho,q)$ where $\rho$ is not the sign irrep, check that the coefficients $c_{i}$ across all neurons corresponding to $(\rho,q)$ is constant. This takes time $O(m)$.
  4. 4.
    For neurons corresponding to the sign irrep, there is only one $\rho$-set $\mathcal{B}_{\rho}=\{\pm 1\}$. Check that the constraint in Observation B.2(8) holds; that is, that the neurons corresponding to $(+1,+1)$ have coefficients summing to the same value as those corresponding to $(-1,-1)$, and likewise for $(+1,-1)$ and $(-1,+1)$.
  5. 5.
    Construct the idealized parameters $\tilde{\bm{\theta}}$ consisting of
    $\tilde{\bm{w}}_{l}^{i}(x)$ $={\bm{b}}_{i}^{\top}\rho(x){\bm{a}}_{i}$
    $\tilde{\bm{w}}_{r}^{i}(y)$ $={\bm{a}}_{i}^{\top}\rho(x){\bm{c}}_{i}$
    $\tilde{\bm{w}}_{u}^{i}(z)$ $=c_{i}{\bm{b}}_{i}^{\top}\rho(z){\bm{b}}_{j}$
    $\tilde{\bm{w}}_{b}(z)$ $=(c_{-}-c_{+})\operatorname{sgn}(z).$
    This takes time $O(\sum_{i=1}^{m}\dim(\rho_{i})^{2}\lvert G\rvert\leq O(\lvert G\rvert^{2}m)$.
  6. 6.
    Compute the idealized margin
    $M=\min_{x,y\in G}f_{\tilde{\bm{\theta}}}(xy\mid x,y)-\max_{z^{\prime}\neq xy}f_{\tilde{\bm{\theta}}}(z\mid xy).$
    By Proposition B.3, this can be done with a single forward pass of $f_{\tilde{\bm{\theta}}}$, in time $O(\lvert G\rvert m)$.
  7. 7.
    Use Lemma F.2 to compute $L$ as in Eq 11 in time $O(m\lvert G\rvert^{2})$. The final accuracy bound is
    $\Pr_{x,y\sim\operatorname{Unif}(G)}[M>L(x,y)].$
The total time complexity is dominated by the last step. Soundness, i.e. $\forall\pi:V_{\mathrm{irrep}}({\bm{\theta}},G,\pi)\leq\operatorname{\alpha_{G}}({\bm{\theta}})$, is obvious from construction and Proposition B.3.
Appendix E Insufficiency of causal interventions
Stander et al. (2024) perform a series of causal interventions on a model trained on $S_{5}$, and find results consistent with their description of the cosets algorithm. However, we expect that these interventions would have the same result for models implementing different algorithms. To verify this, we replicate their outcomes with a model trained on the cyclic group $G={\mathbb{Z}}/53{\mathbb{Z}}$; such a model cannot be using the coset algorithm, as $G$ as no non-trivial subgroups. Models trained cyclic groups were studied in Yip et al. (2024) and found to be using a distinct algorithm; see also discussion in Section B.1.
In detail, the interventions that Stander et al. (2024) perform on $S_{5}$ are:
  1. 1.
    Embedding exchange: Swapping the model’s left and right embeddings destroys model performance. Since $S_{5}$ is non-commutative, we expect this to be the case regardless of what algorithm the model is implementing. Even with ${\mathbb{Z}}/53{\mathbb{Z}}$, which is commutative, we get this result, since ${\bm{W}}_{l}{\bm{E}}_{l}(x)+{\bm{W}}_{r}{\bm{E}}_{r}(y)\neq{\bm{W}}_{l}{\bm{E}}_{r}(y)+{\bm{W}}_{r}{\bm{E}}_{l}(x)$.
  2. 2.
    Switch permutation sign Multiplying either the left or right embeddings individually by $-1$ destroys model performance, while multiplying both preserves model performance. We find this to be the case with ${\mathbb{Z}}/53{\mathbb{Z}}$ as well.
  3. 3.
    Absolute value non-linearity Replacing the ReLU nonlinearity with an absolute value improves model performance. Again, this is the case with ${\mathbb{Z}}/53{\mathbb{Z}}$ as well. We explain this by decomposing $\operatorname{ReLU}(x)=(x+\lvert x\rvert)/2$. By inspecting the summation in Eq 4, we see that the $x/2$ component sums to zero, so only the $\lvert x\rvert/2$ term contributes. Thus, replacing the $\operatorname{ReLU}$ with an absolute value is approximately equivalent to doubling the activations, which reduces loss assuming the model already has near-perfect accuracy.
  4. 4.
    Distribution change Perturbing model activations by $\mathcal{N}(1,1)$ reduces performance to a greater extent than perturbing with $\mathcal{N}(1,-1)$. Again, we observe this with ${\mathbb{Z}}/53{\mathbb{Z}}$.
See Table 1 for results.
Table 1: Causal interventions aggregated over 128 runs on ${\mathbb{Z}}/53{\mathbb{Z}}$ (ours) juxtaposed with the same interventions aggregated over 128 runs on $S_{5}$ (Stander et al., 2024). We train our models with fewer iterations than Stander et al. (2024), resulting in higher base loss. However, the directional effect of each intervention is the same, even though the coset concentration explanation does not hold for ${\mathbb{Z}}/53{\mathbb{Z}}$.
${\mathbb{Z}}/53{\mathbb{Z}}$ (ours) $S_{5}$ (Stander et al., 2024)
Intervention Mean accuracy Mean loss Mean accuracy Mean loss
Base model 99.55% 0.0711 99.99% 1.97e-6
Embedding swap 03.85% 4.15 1% 4.76
Switch left and right sign 99.69% 0.0663 100% 1.97e-6
Switch left sign 00.00% 17.2 0% 22.39
Switch right sign 00.00% 17.2 0% 22.36
Absolute value nonlinearity 99.88% 0.0045 100% 3.69e-13
Perturb $\mathcal{N}(0,1)$ 76.90% 0.829 97.8% 0.0017
[…]
Perturb $\mathcal{N}(1,1)$ 49.80% 1.79 88% 0.029
Perturb $\mathcal{N}(1,-1)$ 83.17% 0.780 98% 0.0021
Appendix F Additional proofs
Lemma F.1.
Let
${\bm{\theta}}$ $=({\bm{U}},{\bm{E}}_{l},{\bm{E}}_{r},{\bm{w}}_{b})=(({\bm{w}}_{u}^{i},{\bm{w}}_{l}^{i},{\bm{w}}_{r}^{i})_{i=1}^{m},{\bm{w}}_{b}),$
$\tilde{\bm{\theta}}$ $=(\tilde{\bm{U}},\tilde{\bm{E}}_{l},\tilde{\bm{E}}_{r},\tilde{\bm{w}}_{b})=((\tilde{\bm{w}}_{u}^{i},\tilde{\bm{w}}_{l}^{i},\tilde{\bm{w}}_{r}^{i})_{i=1}^{m},\tilde{\bm{w}}_{b})$
Then, for any $x,y\in G$,
$\max_{z\in G}$ $\lvert f_{\bm{\theta}}(z\mid x,y)-f_{{\bm{\theta}}^{\prime}}(z\mid x,y)\rvert$
$=\max_{z\in G}\left\lvert\sum_{i=1}^{m}\left({\bm{w}}^{i}_{u}(z)\operatorname{ReLU}[{\bm{w}}^{i}_{l}(x)+{\bm{w}}^{i}_{l}(y)]+{\bm{w}}_{b}(z)-\tilde{\bm{w}}_{u}(z)\operatorname{ReLU}[\tilde{\bm{w}}_{l}(x)+\tilde{\bm{w}}_{l}(y)]-\tilde{\bm{w}}_{b}(z)\right)\right\rvert$
$\leq\max_{z\in G}\left\lvert\sum_{i=1}^{m}({\bm{w}}^{i}_{u}(z)-\tilde{\bm{w}}^{i}_{u}(z))\operatorname{ReLU}[{\bm{w}}^{i}_{l}(x)+{\bm{w}}^{i}_{l}(y)]\right\rvert$
$\qquad+\max_{z\in G}\left\lvert\sum_{i=1}^{m}\tilde{\bm{w}}^{i}_{u}(z)\left(\operatorname{ReLU}[{\bm{w}}^{i}_{l}(x)+{\bm{w}}^{i}_{l}(y)]-\operatorname{ReLU}[\tilde{\bm{w}}^{i}_{l}(x)+\tilde{\bm{w}}^{i}_{l}(y)]\right)\right\rvert$
$\qquad+\max_{z\in G}\lvert{\bm{w}}_{b}(z)-\tilde{\bm{w}}_{b}(z)\rvert$
$=\lVert({\bm{U}}-\tilde{\bm{U}})^{\top}\operatorname{ReLU}[{\bm{E}}_{l}(x)+{\bm{E}}_{r}(y)]\rVert_{\infty}$
$\qquad+\left\lVert\tilde{\bm{U}}^{\top}\left(\operatorname{ReLU}[{\bm{E}}_{l}(x)+{\bm{E}}_{r}(y)]-\operatorname{ReLU}[\tilde{\bm{E}}_{l}(x)+\tilde{\bm{E}}_{r}(y)]\right)\right\rVert_{\infty}$
$\qquad+\lVert{\bm{w}}_{b}-\tilde{\bm{w}}_{b}\rVert_{\infty}$
$\leq\lVert({\bm{U}}-\tilde{\bm{U}})^{\top}\rVert_{2,\infty}\lVert\operatorname{ReLU}[{\bm{E}}_{l}(x)+{\bm{E}}_{r}(y)]\rVert_{2}$
$\qquad+\lVert\tilde{\bm{U}}^{\top}\rVert_{2,\infty}\lVert\left(\operatorname{ReLU}[{\bm{E}}_{l}(x)+{\bm{E}}_{r}(y)]-\operatorname{ReLU}[\tilde{\bm{E}}_{l}(x)+\tilde{\bm{E}}_{r}(y)]\right)\rVert_{2}$
$\qquad+\lVert{\bm{w}}_{b}-\tilde{\bm{w}}_{b}\rVert_{\infty}$
$\leq\lVert({\bm{U}}-\tilde{\bm{U}})^{\top}\rVert_{2,\infty}(\lVert{\bm{E}}_{l}(x)\rVert_{1,2}+\lVert{\bm{E}}_{r}(y)\rVert_{1,2})$
$\qquad+\lVert\tilde{\bm{U}}^{\top}\rVert_{2,\infty}(\lVert{\bm{E}}_{l}-\tilde{\bm{E}}_{l}\rVert_{1,2}+\lVert{\bm{E}}_{r}-\tilde{\bm{E}}_{r}\rVert_{1,2})$
$\qquad+\lVert{\bm{w}}_{b}-\tilde{\bm{w}}_{b}\rVert_{\infty}.$
Lemma F.2.
Let ${\bm{\theta}}$ and $\tilde{\bm{\theta}}$ be as in Lemma F.1. Further, suppose we have a margin lower bound function for ${\bm{\theta}}$
$M(z)\leq\min_{\begin{subarray}{c}x,y,z^{\prime}\in G\\ xy=z\neq z^{\prime}\end{subarray}}f_{\bm{\theta}}(xy\mid x,y)-f_{\bm{\theta}}(z^{\prime}\mid x,y).$
Then,
$\operatorname{\alpha_{G}}(\tilde{\bm{\theta}})\geq\Pr_{x,y\sim\operatorname{Unif}(G)}$ $\bigg{[}M(xy)>\lVert({\bm{U}}-\tilde{\bm{U}})^{\top}\rVert_{2,\infty}\lVert\operatorname{ReLU}[{\bm{E}}_{l}(x)+{\bm{E}}_{r}(y)]\rVert_{2}$ (12)
$\qquad+\lVert\tilde{\bm{U}}^{\top}\rVert_{2,\infty}\left\lVert\left(\operatorname{ReLU}[{\bm{E}}_{l}(x)+{\bm{E}}_{r}(y)]-\operatorname{ReLU}[\tilde{\bm{E}}_{l}(x)+\tilde{\bm{E}}_{r}(y)]\right)\right\rVert_{2}$
$\qquad+\lVert{\bm{w}}_{b}-\tilde{\bm{w}}_{b}\rVert_{\infty}\bigg{]}$
$\geq\Pr_{z\sim\operatorname{Unif}(G)}$ $\bigg{[}M(z)>\lVert({\bm{U}}-\tilde{\bm{U}})^{\top}\rVert_{2,\infty}(\lVert{\bm{E}}_{l}(x)\rVert_{1,2}+\lVert{\bm{E}}_{r}(y)\rVert_{1,2})$ (13)
$\qquad+\lVert\tilde{\bm{U}}^{\top}\rVert_{2,\infty}(\lVert{\bm{E}}_{l}-\tilde{\bm{E}}_{l}\rVert_{1,2}+\lVert{\bm{E}}_{r}-\tilde{\bm{E}}_{r}\rVert_{1,2})$
$\qquad+\lVert{\bm{w}}_{b}-\tilde{\bm{w}}_{b}\rVert_{\infty}\bigg{]}.$
The bound in Equation 12 can be computed in time $O(m\lvert G\rvert^{2})$, while that in Equation 13 can be computed in time $O(m\lvert G\rvert)$.
Proof.
This follows immediately from Lemma F.1. ∎
Lemma F.3.
Let $\rho\colon G\to\operatorname{GL}(n,{\mathbb{R}})$ be a permutation representation of $G$ that decomposes into two subspaces $V\oplus W$ such that $\rho$ acts trivially on $W$. (For example, the standard $(n-1)$-dimensional irrep of $S_{n}$ is of this form.) Let ${\bm{V}}$ and ${\bm{W}}$ be orthonormal bases of $V,W\in{\mathbb{R}}^{n}$, respectively, and let ${\bm{P}}={\bm{V}}{\bm{V}}^{\top}$ and ${\bm{Q}}={\bm{W}}{\bm{W}}^{\top}$ be the corresponding orthogonal projections, so that ${\bm{P}}+{\bm{Q}}={\bm{I}}$. Denote ${\bm{b}}_{i}={\bm{V}}^{\top}{\bm{e}}_{i}$, where $({\bm{e}}_{i})_{i=1}^{n}$ are the standard basis of ${\mathbb{R}}^{n}$. Then, for any ${\bm{a}}\in V$,
$\varphi(z)=\left\langle\rho|_{V}(z),-\sum_{i,j}^{n}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{b}}_{j}{\bm{b}}_{i}^{\top}\right\rangle$
is maximized at $z=e$.
Proof.
$\varphi(z)$ $=-\left\langle\rho|_{V}(z),\sum_{i,j}^{n}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{b}}_{j}{\bm{b}}_{i}^{\top}\right\rangle$
$=-\left\langle\rho|_{V}(z),\sum_{i,j}^{n}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{V}}^{\top}{\bm{e}}_{j}{\bm{e}}_{i}^{\top}{\bm{V}}\right\rangle$
$=-\left\langle{\bm{V}}\rho|_{V}(z){\bm{V}}^{\top},\sum_{i,j}^{n}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{e}}_{j}{\bm{e}}_{i}^{\top}\right\rangle$
$=-\left\langle{\bm{P}}\rho(z){\bm{P}},\sum_{i,j}^{n}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{e}}_{j}{\bm{e}}_{i}^{\top}\right\rangle$
$=-\left\langle\rho(z),\sum_{i,j}^{n}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{e}}_{j}{\bm{e}}_{i}^{\top}\right\rangle$
$\qquad-\left\langle{\bm{Q}}\rho(z){\bm{Q}}^{\top},\sum_{i,j}^{n}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{e}}_{j}{\bm{e}}_{i}^{\top}\right\rangle$
$=-\left\langle\rho(z),\sum_{i,j}^{n}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{e}}_{j}{\bm{e}}_{i}^{\top}\right\rangle$
$\qquad-\left\langle{\bm{Q}}{\bm{Q}}^{\top},\sum_{i,j}^{n}\operatorname{ReLU}[{\bm{a}}^{\top}({\bm{b}}_{i}-{\bm{b}}_{j})]{\bm{e}}_{j}{\bm{e}}_{i}^{\top}\right\rangle,$
where the last step uses that $\rho|_{W}$ is trivial. The first term of the last line is the negation of a sum over off-diagonal entries of the permutation matrix $\rho(z)$, and thus maximized at $z=e$. The second term does not depend on $z$. ∎
Lemma F.4.
Suppose $f\colon G\to{\mathbb{C}}$ is coset concentrated and irrep sparse; that is, $f$ is constant on the cosets of some $H\leq G$ and there exists $\rho\in\operatorname{Irrep}(G)$ mapping $\rho\colon G\to\operatorname{GL}(d,{\mathbb{C}})$ such that $f$ is a linear combination of the entries of $\rho$. Then, there must exist an embedding $\iota\colon G/H\to{\mathbb{C}}^{d}$ such that the action of $\rho$ on $\iota(G/H)$ is isomorphic to the action of $G$ on $G/H$. Further, there exist ${\bm{a}}\in({\mathbb{C}}^{d})^{*}$ and ${\bm{b}}\in{\mathbb{C}}^{d}$ such that $f(g)=\langle{\bm{a}},\rho(g){\bm{b}}\rangle$.
The converse also holds and is immediate.
Proof.
Let $\{g_{1},\ldots,g_{k}\}$ be representatives for the cosets $G/H$ with $g_{1}=e$, and let $\tilde{\rho}\colon G\to S_{k}\subseteq\operatorname{GL}(\bigoplus_{i=1}^{k}g_{i}{\mathbb{C}})=:V$ be the permutation representation corresponding to the action of $G$ on $G/H$. (That is, $\tilde{\rho}$ is induced by the trivial representation on $H$.) Let $\{{\bm{e}}_{g_{i}}\}_{i=1}^{k}$ denote the basis vectors of $V$. Define ${\bm{a}}\in V^{*}$ by $\langle{\bm{a}},{\bm{e}}_{g_{i}}\rangle=f(g_{i})$ for each $i\in[k]$. Hence $f(g_{i})=\langle{\bm{a}},\tilde{\rho}(g_{i}){\bm{e}}_{g_{1}}\rangle$. Since $f$ is coset concentrated, this same relation holds for all $g\in G$. Now, let $V=\bigoplus_{j=1}^{n}V_{j}$ be the decomposition of $V$ into irreps, with corresponding projections $\pi_{j}\colon V\to V_{j}$ and irreps $\rho_{j}\colon G\to\operatorname{GL}(V_{j})$. We have
$f(g)=\langle{\bm{a}},\tilde{\rho}(g){\bm{e}}_{g_{1}}\rangle=\sum_{j=1}^{n}\langle{\bm{a}}|_{V_{j}},\pi_{j}\tilde{\rho}(g){\bm{e}}_{g_{1}}\rangle=\sum_{j=1}^{n}\langle{\bm{a}}|_{V_{j}},\rho_{j}(g)\pi_{j}{\bm{e}}_{g_{1}}\rangle.$
By the irrep sparsity assumption, at most one term of this sum is nonzero, and that corresponding term must have $\rho_{j}=\rho$:
$f(g)=\langle{\bm{a}}|_{V_{j}},\rho(g)\pi_{j}{\bm{e}}_{g_{1}}\rangle.$
Then $\rho$ acts on $\{\pi_{j}{\bm{e}}_{g_{i}}\}_{i=1}^{k}$ isomorphically to the action of $G$ on $G/H$ and $({\bm{a}}|_{V_{j}},\pi_{j}{\bm{e}}_{g_{1}})$ is the desired covector and vector pair from the theorem statement. ∎
Appendix G Permutation representations and $\rho$-sets of $S_{5}$
Irrep Minimal $\rho$-set size Stabilizer
trivial (1d-0) 1 $S_{5}$
sign (1d-1) 2 $A_{5}$
standard (4d-0) 5 $S_{4}$
sign-standard (4d-1) 10 $A_{4}$
5d-0 6 $F_{5}$
5d-1 12 $D_{12}$
6d-0 20 ${\mathbb{Z}}/6{\mathbb{Z}}$ or $S_{3}$
Table 2: Irreps $\rho\in\operatorname{Irrep}(S_{5})$ by the size of the minimal $\rho$-set, and corresponding stabilizers. We name each irrep by its dimension and an arbitrary disambiguating integer; e.g. 5d-0 is a five-dimensional irrep. A projected $\rho$-set (Definition 7.3) is constant on the cosets of its stabilizer. Notice that the ordering of $\operatorname{Irrep}(S_{5})$ by minimal $\rho$-set size matches the ordering by frequencies with which irreps are learned (Chughtai et al., 2023, Figure 7).
Stabilizer $G$-set size Irreps present
$S_{5}$ 1 1d-0
$A_{5}$ 2 1d-0,1d-1
$S_{4}$ 5 1d-0,4d-0
$F_{5}$ 6 1d-0,5d-0
$A_{4}$ 10 1d-0,1d-1,4d-0,4d-1
$D_{12}$ 10 1d-0,4d-0,5d-1
$D_{10}$ 12 1d-0,1d-1,5d-0,5d-1
$D_{8}$ 15 1d-0,4d-0,5d-0,5d-1
${\mathbb{Z}}/6{\mathbb{Z}}$ 20 1d-0,4d-0,4d-1,5d-1,6d-0
$S_{3}^{0}$ 20 1d-0,4d-0,5d-1,6d-0
$S_{3}^{1}$ 20 1d-0,1d-1,4d-0,4d-1,5d-0,5d-1
${\mathbb{Z}}/5{\mathbb{Z}}$ 24 1d-0,1d-1,5d-0,5d-1,6d-0
$V_{4}^{0}$ 30 1d-0,1d-1,4d-0,4d-1,5d-0,5d-1
$V_{4}^{1}$ 30 1d-0,4d-0,5d-0,5d-1,6d-0
${\mathbb{Z}}/4{\mathbb{Z}}$ 30 1d-0,4d-0,4d-1,5d-0,5d-1,6d-0
Table 3: All transitive permutation representations of $S_{5}$ with size no more than 30 along with decomposition into irreps. By the orbit-stabilizer theorem, transitive permutation representations correspond directly to left actions of $S_{5}$ on cosets of its subgroups; the subgroup acted upon is then the stabilizer of the action. Upper indices disambiguate subgroups of $S_{5}$ that are isomorphic but not conjugate. $F_{5}$ is the Frobenius group of order $20$ and $V_{4}$ is the Klein four-group.
Appendix H Experiment details
We train 100 training runs of one-hidden-layer models with hidden dimensionality $m=128$ on the group $S_{5}$. The test set is all pairs of two inputs from $S_{5}$, with $\lvert S_{5}\rvert^{2}=14400$ points total. The training set comprises iid samples from the test set and has size 40% of the test set. Note that we use the same training set for each of the 100 training runs. Each model was trained over 25000 epochs. Learning rate was set to 1e-2. We use the Adam optimizer (Kingma & Ba, 2015) with weight decay 2e-4 and $(\beta_{1},\beta_{2})=(0.9,0.98)$. [^18] Note that previous work uses AdamW (Loshchilov & Hutter, 2019) instead, with weight decay 1. However, we found that models trained with Adam grok the group composition task in an order of magnitude fewer epochs. We did not notice significant differences in the end result post-grokking between models trained with Adam vs AdamW. All models were trained on one Nvidia A6000 GPU. Compact proof verifiers were run on an Intel Core i5-1350P CPU. Neural networks were implemented in PyTorch (Paszke et al., 2019). Their group-theoretic properties were analyzed with GAP (GAP, 2024).
As input to the verifier $V_{\mathrm{coset}}$, recall that the interpretation string looks like $\pi=((H_{i},g_{i}))_{i=1}^{m}$, where the left embedding at neuron $i\in[m]$ should be constant on the right cosets of $H_{i}$ and the right embedding at neuron $i$ should be constant on the left cosets of $g_{i}H_{i}g_{i}^{-1}$. When constructing $\pi$, we set each $H_{i}$ to be the largest subgroup of $G$ such that
$\dfrac{\mathbb{E}\operatorname{Var}({\bm{w}}_{l}^{i}\mid H_{i}\backslash G)}{\operatorname{Var}({\bm{w}}_{l}^{i})}<0.01,$
and $K_{i}$ to the largest subgroup such that
$\dfrac{\mathbb{E}\operatorname{Var}({\bm{w}}_{l}^{i}\mid G/K_{i})}{\operatorname{Var}({\bm{w}}_{l}^{i})}<0.01,$
and check for the existence of $g_{i}$ such that $K_{i}=g_{i}H_{i}g_{i}^{-1}$. The quotient on the LHS is bounded in $[0,1]$ by the law of total variance.
As input to the verifier $V_{\mathrm{irrep}}$, the interpretation string labels each neuron with an $\rho_{i}\in\operatorname{Irrep}(G)$ and a corresponding $\rho_{i}$-set $\mathcal{B}$. The irrep $\rho_{i}$ is chosen to have the largest $R^{2}$ against ${\bm{w}}_{l}^{i}$, or none, if the $R^{2}$ if no irrep exceeds $95\%$. The $\rho$-set is recovered using singular value decomposition of the coefficient matrices ${\bm{A}}_{i},{\bm{B}}_{i},{\bm{C}}_{i}$ as defined in Observation B.2.
Recall that the construction of the interpretation string $\pi$ does not count towards the runtime of the verifier.
Appendix I Beyond permutation groups
In the main text, we focus on the group $S_{5}$. This group has only real irreps, in the sense that every irrep over ${\mathbb{C}}$ is isomorphic to an irrep with only real matrix entries; equivalently, every irrep has positive Frobenius-Schur indicator $\iota(\rho):=\lvert G\rvert^{-1}\sum_{g\in G}\operatorname{tr}(\rho(g))$. In general, irreps have Frobenius-Schur indicator in $\{1,0,-1\}$, corresponding to the irrep being real, complex, and quaternionic respectively. These three cases correspond to the ring of $G$-linear endomorphisms of the irrep being isomorphic to either ${\mathbb{R}},{\mathbb{C}},{\mathbb{Q}}$. By Schur’s lemma, the endomorphism ring is a real associative division algebra, so these are the only three cases.
In preliminary investigations of more general irreps $\rho$, we convert irreps over ${\mathbb{C}}$ with nonpositive Frobenius-Schur indicator to irreps over ${\mathbb{R}}$ of twice the dimensionality:
$\tilde{\rho}(g)=\begin{bmatrix}\operatorname{Re}\rho(g)&-\operatorname{Im}\rho(g)\\ \operatorname{Im}\rho(g)&\operatorname{Re}\rho(g).\end{bmatrix}$ (14)
We then find that the ${\bm{A}},{\bm{B}},{\bm{C}}$ matrices of Observation B.2(1) are approximately rank one when $\rho$ is real, rank two when $\rho$ is complex, and rank four when $\rho$ is quaternionic. In the complex case, we find also that when ${\mathbb{R}}^{d}$ is given the complex structure induced by Eq 14, the two singular vectors are conjugate, and correspond to equal singular values. Thus, we speculate that the neural network uses the same $\rho$-sets algorithm as in the real case, but over ${\mathbb{C}}$, and then takes the real part: letting $\rho\in\operatorname{GL}(n,{\mathbb{C}})$ and $\mathcal{B}\subseteq{\mathbb{C}}^{d}$ a finite $\rho$-set,
$f_{\rho,\mathcal{B}}(z\mid x,y)=-\sum_{{\bm{b}},{\bm{b}}^{\prime}\in\mathcal{B}}\operatorname{Re}{\bm{b}}^{\top}\rho(z){\bm{b}}^{\prime}\operatorname{ReLU}[\operatorname{Re}({\bm{b}}^{\top}\rho(x){\bm{a}}-{\bm{a}}^{\top}\rho(y){\bm{b}}^{\prime})].$